MASS-ANALYZED THRESHOLD IONIZATION OF LaO_{2}

\underline{LU WU}, CHANGHUA ZHANG, SERGIY KRASNOKUTSKI, and DONG-SHENG YANG, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

Lanthanum oxide, LaO_{2}, is produced in a pulsed laser-vaporization metal-cluster source and studied by mass-analyzed threshold ionization (MATI) spectroscopy. From the MATI spectrum, the adiabatic ionization energy of LaO_{2} is determined to be 40134 (5) cm^{−}^{1} or 4.976 (6) eV, and La^{+}-O stretching and O-La^{+}-O bending frequencies are measured as 656 and 120 cm^{−}^{1}. The measured ionization energy is about 3.0 eV lower than the value predicted by recent high-level ab initio calculations.^{a} In this talk, we will discuss the discrepancy between the experiment and theory and the electronic transition observed in our experiment.

aT. K. Todorova, I. Infante, L. Gagliardi, and J. M. Dyke, J. Phys. Chem. A 112, 7825 (2008).